OpenStudy (anonymous):

Need answer and how its done. (1/x-2)+(2x/(x-2)(x-8))=(x/2(x-8))

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ x-2\end{matrix}\right) + \left(\begin{matrix}2x \\ (x-2)(x-8)\end{matrix}\right) = \left(\begin{matrix}x \\ 2(x-8)\end{matrix}\right)\]

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

can you show work? @quietlysinging

OpenStudy (zzr0ck3r):

@Chrisgoblin use `\(\frac{a}{b}\)` for \(\frac{a}{b}\)

OpenStudy (anonymous):

what?

OpenStudy (zzr0ck3r):

for fractions.

OpenStudy (anonymous):

can you just show me how its done please I will fan

OpenStudy (zzr0ck3r):

`\(\frac{1}{x-2}+\frac{2x}{(x-2)(x-8)}=\frac{x}{x(x-8)}\) ` \(\huge\downarrow\) \(\frac{1}{x-2}+\frac{2x}{(x-2)(x-8)}=\frac{x}{x(x-8)}\)

OpenStudy (zzr0ck3r):

you dont need to fan me...

OpenStudy (anonymous):

what does \(\frac{1}{x-2}+\frac{2x}{(x-2)(x-8)}=\frac{x}{x(x-8)}\) mean

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

thats the question

OpenStudy (zzr0ck3r):

I am just showing you how to make this \[\left(\begin{matrix}1 \\ x-2\end{matrix}\right) + \left(\begin{matrix}2x \\ (x-2)(x-8)\end{matrix}\right) = \left(\begin{matrix}x \\ 2(x-8)\end{matrix}\right)\] look like this \[\frac{1}{x-2}+\frac{2x}{(x-2)(x-8)}=\frac{x}{x(x-8)}\]

OpenStudy (anonymous):

ok do you know how to answer it though?

OpenStudy (zzr0ck3r):

oh sure \(\frac{1}{x-2}+\frac{2x}{(x-2)(x-8)}=\frac{x}{2(x-8)}\) multiply every term by \(2(x-2)(x-8)\) and you get \(2(x-8)+2*2x=x(x-2)\iff 2x-16+4x=x^2-2x \) so \(x^2-8x+16=0\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that the answe? z^2-8x+16=0? if it is thanks

OpenStudy (zzr0ck3r):

well the answer is \((x-4)(x-4)=0 \implies x=4\)